If the perimeter of a right triangle is 60 and the sides are x, 2x + 4, and 3x – 4, what is the area of the triangle?
Answer by mananth(16946)

x +2x + 4+ 3x – 4 =60
6x=60
x=10
..
2x+4 = 2*10+4 =24
..
3x-4 = 3*10-4=26
the right triangle
10,24,26
..
Area = 1/2 * 10*24
=120 sq.units